# Verina: Benchmarking Verifiable Code Generation

### v1

@@ A.1 License
We ensure compliance with all relevant licenses: MBPP-DFY-50 [18] is licensed under GPL-3.0, while both CloverBench [12] and LiveCodeBench [1] use MIT licenses. Our datasets Verina is licensed under GPL-3.0. Consistent with established research practices [48, 1], we only use publicly available materials from competitive programming platforms such as LeetCode. Our collection and use of these problems is strictly for academic research purposes, and Verina involves no model training or fine-tuning processes.
@@ A.2 Model Configurations and Compute
Table 3 presents the configuration details and total experiment costs for all nine evaluated LLMs. For all LLMs, we use a temperature of 1.0 and a maximum output token budget of 10,000. For reasoning models, we use default settings of reasoning efforts or budgets. We host DeepSeek Prover 2 7B and Qwen 3 235B-A22B locally using 8 NVIDIA H100 80GB GPUs. We run other LLMs through APIs, for which we provide the total cost and cost per million tokens. The costs marked with asterisks include the additional expenses incurred during iterative proof refinement experiments, which required up to 64 refinement attempts per datapoint.
@@
DeepSeek Prover 2 7B DeepSeek-Prover-V2-7B GPU - -
@@ A.3 Prompts
We employ a consistent 2-shot prompting approach across all models and tasks to enhance output format adherence and task understanding. The 2-shot examples are excluded from the final benchmark evaluation. For each problem instance, we sample 5 responses from each model and calculate pass@1 metrics [3] using these 5 samples to ensure robust evaluation statistics. We utilize DSPy [49] for structural prompting. We provide the detailed prompts in the following: Prompt A.3 for CodeGen, Prompt A.3 for SpecGen, Prompt A.3 for ProofGen, and Prompt A.3 for ProofGen with iterative refinement.
{prompt}
[CodeGen ] Instructions
You are an expert in Lean 4 programming and theorem proving.
Please generate a Lean 4 program that finishes the task described in
task_description using the template provided in task_template‘.
The task_template is a Lean 4 code snippet that contains placeholders
(warpped with {{}}) for the code to be generated.
The program should:
- Be well-documented with comments if necessary
- Follow Lean 4 best practices and use appropriate Lean 4 syntax and features
- DO NOT use Lean 3 syntax or features
- DO NOT import Std or Init
Hint:
- Use a[i]! instead of a[i] when a is an array or a list when necessary
Input Fields
  • task_description
@@
    Lean 4 template with placeholders for code generation and optional reference
    specification.
Output Fields
  • imports
@@
    Generated Lean 4 code following the template signature and complete the task.
{prompt}
[SpecGen ] Instructions
You are an expert in Lean 4 programming and theorem proving.
Please generate a Lean 4 specification that constrains the program
implementation using the template provided in task_template‘.
The task_template is a Lean 4 code snippet that contains placeholders
(warpped with {{}}) for the spec to be generated.
The precondition should be as permissive as possible, and the postcondition
should model a sound an complete relationship between input and output of the
program based on the task_description‘.
The generated specification should:
- Be well-documented with comments if necessary
- Follow Lean 4 best practices and use appropriate Lean 4 syntax and features
- DO NOT use Lean 3 syntax or features
- DO NOT import Std or Init
- Only use precond_aux or postcond_aux when you cannot express
the precondition or postcondition in the main body of the specification
- add @[reducible, simp] attribute to the definitions in precond_aux or
postcond_aux
Hint:
- Use a[i]! instead of a[i] when a is an array or a list when necessary
Input Fields
  • task_description
@@
    Lean 4 template with placeholders for specfication generation and optional
    reference code.
Output Fields
  • imports
    Imports needed for precond and postcond‘. Keep it empty if not needed.
@@
{prompt}
[ProofGen ] Instructions
You are an expert in Lean 4 programming and theorem proving.
Please generate a Lean 4 proof that the program satisfies the specification
using the template provided in task_template‘.
The task_template is a Lean 4 code snippet that contains placeholders
(warpped with {{}}) for the proof to be generated.
The proof should:
- Be well-documented with comments if necessary
- Follow Lean 4 best practices and use appropriate Lean 4 syntax and features
- DO NOT use Lean 3 syntax or features
- DO NOT import Std or Init
- DO NOT use cheat codes like sorry
Hint:
- Unfold the implementation and specification definitions when necessary
- Unfold the precondition definitions at h_precond when necessary
Input Fields
  • task_description
@@
    Lean 4 template with code and specification to be proved, and placeholders
    for proof generation.
Output Fields
  • imports
@@
    Auxiliary definitions and lemma for proof‘. Keep it empty if not needed.
@@
{prompt}
[ProofGen with Iterative Refinement] Instructions
You are an expert in Lean 4 programming and theorem proving.
Please generate a Lean 4 proof that the program satisfies the specification
using the template provided in task_template‘.
The task_template is a Lean 4 code snippet that contains placeholders
(warpped with {{}}) for the proof to be generated.
The proof should:
- Be well-documented with comments if necessary
- Follow Lean 4 best practices and use appropriate Lean 4 syntax and features
- DO NOT use Lean 3 syntax or features
- DO NOT import Std or Init
- DO NOT use cheat codes like sorry
Hint:
- Unfold the implementation and specification definitions when necessary
- Unfold the precondition definitions at h_precond when necessary
Furthermore, prev_error is the error message from the previous proving
attempt.
Please use the prev_imports‘, prev_proof_aux‘, and prev_proof as
references to improve the generated proof.
- You can ignore unused variable warnings in the error message.
Input Fields
  • task_description
@@
    Lean 4 template with code and specification to be proved, and placeholders
    for proof generation.
@@
Output Fields
  • imports
@@
    Auxiliary definitions and lemma for proof‘. Keep it empty if not needed.
@@
A.4 Implementation of Evaluation Metrics in Lean
@@
Specification evaluation. Recall in Section 4.1, we define the soundness and completeness of model-generated pre-condition $\hat{P}$ and post-condition $\hat{Q}$ in relation to their ground truth counterparts $P$ and $Q$: (i) $\hat{P}$ is sound iff $\forall\overline{x}{}.P{}(\overline{x}{}){}\Rightarrow\hat{P}{}(\overline{x}{}){}$; (ii) $\hat{P}$ is complete iff $\forall\overline{x}{}.\hat{P}{}(\overline{x}{}){}\Rightarrow P{}(\overline{x}{}){}$; (iii) $\hat{Q}$ is sound iff $\forall\overline{x}{},y{}.P{}(\overline{x}{}){}\,\land\,\hat{Q}{}(\overline{x}{},y{}){}\Rightarrow Q{}(\overline{x}{},y{}){}$; (iv) $\hat{Q}$ is complete iff $\forall\overline{x}{},y{}.P{}(\overline{x}{}){}\,\land\,Q{}(\overline{x}{},y{}){}\Rightarrow\hat{Q}{}(\overline{x}{},y{}){}$. Moreover, since our evaluator is based on testing, we only require that $\overline{x}$ and $y$ are from our test suite. Our quality assurance process in Section 3.2 ensures that all ground truth pre-conditions and post-conditions pass our positive tests and do not pass our negative tests. Therefore, we can simplify the soundness and completeness metrics as follows:
@@ Appendix B Additional Experimental Evaluation Results
Detailed performance breakdown. Tables 4, 5 and 6 provide detailed breakdowns of model performance across the three foundational tasks. They reveal that syntax incorrectness and use of non-existent library functions (as demonstrated in Appendix C) represent the major problems, especially for less capable models. Specifically, after manual inspection of the evaluation result, DeepSeek Prover 2 7B and Qwen 3 235B-A22B-FP8 suffer from instruction following ability, failing to output the desired format specified in our prompts (cf. Section A.3). The relatively low unknown percentages across most evaluations demonstrate that our specification evaluation metric is reliable. Pre-conditions are generally simpler than post-conditions, resulting in lower unknown rates during evaluation. More capable models often generate specifications with more complicated logical structures, leading to higher unknown percentages in post-condition evaluation. We present a case study in Appendix C on the challenge of automatically evaluating LLM-generated specifications. In our main results, we report the uncertainty from unknown cases using error bars, where the lower bound represents the Passin the table and the upper bound represents Pass%+Unknownin the table.
@@
Model Cannot CompileFail Unit TestPass%
@@
Model Cannot CompileSoundness Completeness
PassFailUnknownPassFailUnknown%
@@
Model Cannot CompileSoundness Completeness
PassFailUnknownPassFailUnknown%
@@ Appendix C Case Studies of Model Failures and Evaluation Metrics
Figure 13: Example (verina_advanced_1): o4-mini generates code with hallucinated Int.xor method that does not exist in Lean 4.
@@
Figure 14: Example (verina_advanced_57): o4-mini generates code with logical error in state management, causing unit test failure.
@@
Figure 15: Example (verina_advanced_76): o4-mini generates unsound pre-condition using strict inequality instead of allowing k to equal the number of distinct elements.
Specification generation failures: incomplete pre-conditions. Figure 16 demonstrates how LLMs can generate overly permissive preconditions that fail to capture essential constraints. The task description specifies that “All integers in both arrays are unique" and that "nums1: A list of integers, which is a subset of nums2”. The ground truth precondition correctly enforces three critical requirements: List.Nodup nums1 ensures uniqueness in the first array, List.Nodup nums2 ensures uniqueness in the second array, and nums1.all (fun x => x nums2) verifies that nums1 is indeed a subset of nums2. However, the LLM-generated precondition simply uses True, completely ignoring all stated constraints. This makes the precondition incomplete as it accepts invalid inputs that violate the problem’s fundamental assumptions, potentially leading to incorrect behavior in the implementation and proof generation phases. In our test suites, we have a negative test case with nums1 = [1, 1] and nums2 = [1, 2]. Since the LLM-generated pre-condition accepts this negative test case, our evaluation metric determines that the LLM-generated pre-condition is incomplete.

### v2

@@ A.1 License
We ensure compliance with all relevant licenses: MBPP-DFY-50 (Misu et al., 2024) is licensed under GPL-3.0, while both CloverBench (Sun et al., 2024) and LiveCodeBench (Jain et al., 2025) use MIT licenses. Our datasets Verina will be licensed under GPL-3.0. Consistent with established research practices (Hendrycks et al., 2021; Jain et al., 2025), we only use publicly available materials from competitive programming platforms such as LeetCode. Our collection and use of these problems is strictly for academic research purposes, and Verina involves no model training or fine-tuning processes.
@@ A.2 Model Configurations and Compute
Table 3 presents the configuration details and total experiment costs for all ten evaluated LLMs. For all LLMs, we use a temperature of 1.0 and a maximum output token budget of 10,000. For reasoning models, we use default settings of reasoning efforts or budgets. We host DeepSeek Prover V2 7B, Goedel Prover V2 32B, and Qwen 3 235B-A22B locally using 8 NVIDIA H100 80GB GPUs. We run other LLMs through APIs, for which we provide the total cost and cost per million tokens. The costs marked with asterisks include the additional expenses incurred during iterative proof refinement experiments, which required up to 64 refinement attempts per datapoint.
@@
DeepSeek Prover V2 7B DeepSeek-Prover-V2-7B GPU - -
@@ A.3 Prompts
We employ a consistent 2-shot prompting approach across all models and tasks to enhance output format adherence and task understanding. The 2-shot examples are excluded from the final benchmark evaluation. For each problem instance, we sample 5 responses from each model and calculate pass@1 metrics (Chen et al., 2021) using these 5 samples to ensure robust evaluation statistics. We utilize DSPy (Khattab et al., 2024) for structural prompting. We provide the detailed prompts in the following: Prompt A.3 for CodeGen, Prompt A.3 for SpecGen, Prompt A.3 for ProofGen, and Prompt A.3 for ProofGen with iterative refinement. For DeepSeek Prover V2 7B and Goedel Prover V2 32B, we used their own prompt templates for ProofGen to achieve optimal performance.
Prompt 1 (CodeGen) Instructions You are an expert in Lean 4 programming and theorem proving. Please generate a Lean 4 program that finishes the task described in task_description using the template provided in task_template‘. The task_template is a Lean 4 code snippet that contains placeholders (warpped with {{}}) for the code to be generated. The program should: - Be well-documented with comments if necessary - Follow Lean 4 best practices and use appropriate Lean 4 syntax and features - DO NOT use Lean 3 syntax or features - DO NOT import Std or Init Hint: - Use a[i]! instead of a[i] when a is an array or a list when necessary Input Fields task_description
@@
Lean 4 template with placeholders for code generation and optional
reference specification.
Output Fields imports
@@
Generated Lean 4 code following the template signature and complete
the task.
Prompt 2 (SpecGen) Instructions You are an expert in Lean 4 programming and theorem proving. Please generate a Lean 4 specification that constrains the program implementation using the template provided in task_template‘. The task_template is a Lean 4 code snippet that contains placeholders (warpped with {{}}) for the spec to be generated. The precondition should be as permissive as possible, and the postcondition should model a sound an complete relationship between input and output of the program based on the task_description‘. The generated specification should: - Be well-documented with comments if necessary - Follow Lean 4 best practices and use appropriate Lean 4 syntax and features - DO NOT use Lean 3 syntax or features - DO NOT import Std or Init - Only use precond_aux or postcond_aux when you cannot express the precondition or postcondition in the main body of the specification - add @[reducible, simp] attribute to the definitions in precond_aux or postcond_aux Hint: - Use a[i]! instead of a[i] when a is an array or a list when necessary Input Fields task_description
@@
Lean 4 template with placeholders for specfication generation and
optional reference code.
Output Fields imports
Imports needed for precond and postcond‘. Keep it empty if not
needed.
@@
Prompt 3 (ProofGen) Instructions You are an expert in Lean 4 programming and theorem proving. Please generate a Lean 4 proof that the program satisfies the specification using the template provided in task_template‘. The task_template is a Lean 4 code snippet that contains placeholders (warpped with {{}}) for the proof to be generated. The proof should: - Be well-documented with comments if necessary - Follow Lean 4 best practices and use appropriate Lean 4 syntax and features - DO NOT use Lean 3 syntax or features - DO NOT import Std or Init - DO NOT use cheat codes like sorry Hint: - Unfold the implementation and specification definitions when necessary - Unfold the precondition definitions at h_precond when necessary Input Fields task_description
@@
Lean 4 template with code and specification to be proved, and
placeholders for proof generation.
Output Fields imports
@@
Auxiliary definitions and lemma for proof‘. Keep it empty if not
needed.
@@
Prompt 4 (ProofGen with Iterative Refinement) Instructions You are an expert in Lean 4 programming and theorem proving. Please generate a Lean 4 proof that the program satisfies the specification using the template provided in task_template‘. The task_template is a Lean 4 code snippet that contains placeholders (warpped with {{}}) for the proof to be generated. The proof should: - Be well-documented with comments if necessary - Follow Lean 4 best practices and use appropriate Lean 4 syntax and features - DO NOT use Lean 3 syntax or features - DO NOT import Std or Init - DO NOT use cheat codes like sorry Hint: - Unfold the implementation and specification definitions when necessary - Unfold the precondition definitions at h_precond when necessary Furthermore, prev_error is the error message from the previous proving attempt. Please use the prev_imports‘, prev_proof_aux‘, and prev_proof as references to improve the generated proof. - You can ignore unused variable warnings in the error message. Input Fields task_description
@@
Lean 4 template with code and specification to be proved, and
placeholders for proof generation.
@@
Output Fields imports
@@
Auxiliary definitions and lemma for proof‘. Keep it empty if not
needed.
@@
A.4 Comparison with CLEVER
As summarized in Table 4, CLEVER (Thakur et al., 2025) only supports evaluation of specification generation and specification-guided code generation. It lacks evaluation support for code generation, proof generation, specification inference from code, and fully end-to-end verifiable code generation. In contrast, Verina fully covers all three foundational tasks and their flexible combinations, enabling a more comprehensive assessment of realistic verification workflows.
Moreover, CLEVER’s SpecGen evaluation assumes access to a sound and complete ground truth specification for certification. However, if such ground truth specification is already available, there is little practical value in generating another, as developers would simply use the existing one. This reliance on ground truth specifications therefore limits CLEVER’s applicability and prevents it from reflecting real-world scenarios. In contrast, Verina employs a combined evaluation framework for specification (Section 4.1) leveraging both formal proving and comprehensive testing, which can reliably assess specification quality even when formal proofs are inconclusive.
Table 4: A detailed comparison of Verina with the concurrent work CLEVER (Thakur et al., 2025) on supported tasks in verifiable code generation. ● means fully supported, ○ means unsupported.
Foundational Tasks (Section 4.1) Task Combinations (Section 4.2)
CodeGen SpecGen ProofGen Specification-Guided Specification Inference End-to-End
Code Generation From Code Verifiable Code Generation
(Desc $\to$ Code) (Desc $\to$ Spec) (Code+Spec $\to$ Proof) (Desc + Spec $\to$ Code + Proof) (Desc + Code $\to$ Spec + Proof) (Desc $\to$ Code + Spec + Proof)
CLEVER (Thakur et al., 2025)
Verina
A.5 Implementation of Evaluation Metrics in Lean
@@
Specification evaluation. Recall in Section 4.1, we define the soundness and completeness of model-generated pre-condition $\hat{P}$ and post-condition $\hat{Q}$ in relation to their ground truth counterparts $P$ and $Q$: (i) $\hat{P}$ is sound iff $\forall\overline{x}{}.P{}(\overline{x}{}){}\Rightarrow\hat{P}{}(\overline{x}{}){}$; (ii) $\hat{P}$ is complete iff $\forall\overline{x}{}.\hat{P}{}(\overline{x}{}){}\Rightarrow P{}(\overline{x}{}){}$; (iii) $\hat{Q}$ is sound iff $\forall\overline{x}{},y{}.P{}(\overline{x}{}){}\,\land\,\hat{Q}{}(\overline{x}{},y{}){}\Rightarrow Q{}(\overline{x}{},y{}){}$; (iv) $\hat{Q}$ is complete iff $\forall\overline{x}{},y{}.P{}(\overline{x}{}){}\,\land\,Q{}(\overline{x}{},y{}){}\Rightarrow\hat{Q}{}(\overline{x}{},y{}){}$.
Our specification evaluation pipeline first attempts to establish the soundness and completeness of generated specifications against the ground truth using LLM-based provers. When the proving step is inconclusive, the evaluator proceeds to testing, where we only require that $\overline{x}$ and $y$ are from our test suite. Our quality assurance process in Section 3.2 ensures that all ground truth pre-conditions and post-conditions pass our positive tests and do not pass our negative tests. Therefore, we can simplify the soundness and completeness metrics as follows:
@@
To further examine the role of proofs within our evaluation pipeline, we analyze how often LLM-based provers succeed in establishing the soundness and completeness of generated specifications against the ground truth. In this setup, we use o4-mini and Claude Sonnet 3.7 to construct Lean proofs for the required logical relationships and compare the results with the testing-based evaluation results. Table 5 summarizes the outcomes. Proof success rates are very low, below 4% across all cases, while testing recognizes more than 40% of generated specifications as sound and complete. We have examined all specifications marked as sound and complete by formal proofs. We observe that whenever proofs succeed they always agree with testing, confirming their validity. However, when proofs fail but testing reports correctness, manual inspection of 20 randomly selected disagreements shows that the testing outcome is always correct.
These results indicate that while proofs provide the formal guarantees of the evaluation results when they succeed, current LLM provers are incapable of serving as a reliable metric with high inconclusive rates. Testing-based evaluation methods achieve high empirical accuracy and reliably identify sound and complete specifications even when proofs are inconclusive and therefore play an important role in ensuring robust and comprehensive specification evaluation when the proving-based evaluation is inconclusive.
Table 5: Evaluation of generated specifications for soundness and completeness. Rows indicate the model that generated the specification, while columns indicate the prover used to check correctness. The last column shows results from our testing-based evaluation.
Spec generated by Proved sound and complete by (%) Sound and complete by testing (%)
o4-mini Claude Sonnet 3.7
o4-mini 3.7 1.6 51.0
Claude Sonnet 3.7 3.7 2.6 41.6
@@ Appendix B Additional Experimental Evaluation Results
Based on the construction methodology of Verina datasets in Section 3.2, we categorize the problems translated from human-written Dafny datasets as Verina-A and the problems written from scratch as Verina-B.
Verina-B is much more challenging than Verina-A. The comparison between Verina-A and Verina-B in Figure 12 reveals substantial difficulty gaps on all three tasks. This demonstrates that problem complexity significantly impacts all aspects of verifiable code generation, and Verina-B provides a valuable challenge for advancing future research in this domain.
Figure 12: pass@1@1 performance on three foundational tasks for Verina-A and Verina-B.
@@
Naive proof refinement gains diminish when problem is difficult. As shown in Figure 14, iterative proof refinement yields substantial improvements on simpler problems but only modest gains on more complex ones. For example, o4-mini improves from 7.41% to 22.22% on Verina-A after 64 iterations, while on Verina-B the success rate rises only from 1.23% to 6.17%. Specialized provers like Goedel Prover V2 and DeepSeek Prover V2 generally outperform general-purpose models, yet o4-mini remains surprisingly competitive on difficult instances, achieving stronger iterative refinement gains on Verina-B. This suggests that while verifier feedback is crucial, naive refinement strategies struggle to overcome the inherent complexity of challenging proofs, and that general-purpose LLMs can still contribute meaningfully in difficult settings.
Figure 14: Breakdown of iterative refinement versus direct generation across different subsets. Refinement yields large gains on Verina-A but limited improvements on Verina-B.
Detailed performance breakdown. Tables 6, 7 and 8 provide detailed breakdowns of model performance across the three foundational tasks. They reveal that syntax incorrectness and use of non-existent library functions (as demonstrated in Appendix C) represent the major problems, especially for less capable models. Specifically, after manual inspection of the evaluation result, Qwen 3 235B-A22B-FP8 suffers from instruction following ability, failing to output the desired format specified in our prompts (cf. Section A.3). The relatively low unknown percentages across most evaluations demonstrate that our specification evaluation metric is reliable. Pre-conditions are generally simpler than post-conditions, resulting in lower unknown rates during evaluation. More capable models often generate specifications with more complicated logical structures, leading to higher unknown percentages in post-condition evaluation. We present a case study in Appendix C on the challenge of automatically evaluating LLM-generated specifications. In our main results, we report the uncertainty from unknown cases using error bars, where the lower bound represents the Pass% in the table and the upper bound represents Pass%+Unknown% in the table.
@@
Model Cannot Compile% Fail Unit Test% Pass%
@@
Model Cannot Compile% Soundness Completeness
Pass% Fail% Unknown% Pass% Fail% Unknown%
@@
Model Cannot Compile% Soundness Completeness
Pass% Fail% Unknown% Pass% Fail% Unknown%
@@ Appendix C Case Studies of Model Failures and Evaluation Metrics
Figure 15: Example (verina_b_1): o4-mini generates code with hallucinated Int.xor method that does not exist in Lean 4.
@@
Figure 16: Example (verina_b_57): o4-mini generates code with logical error in state management, causing unit test failure.
@@
Figure 17: Example (verina_b_76): o4-mini generates unsound pre-condition using strict inequality instead of allowing k to equal the number of distinct elements.
Specification generation failures: incomplete pre-conditions. Figure 18 demonstrates how LLMs can generate overly permissive preconditions that fail to capture essential constraints. The task description specifies that “All integers in both arrays are unique” and that ”nums1: A list of integers, which is a subset of nums2”. The ground truth precondition correctly enforces three critical requirements: List.Nodup nums1 ensures uniqueness in the first array, List.Nodup nums2 ensures uniqueness in the second array, and nums1.all (fun x => x nums2) verifies that nums1 is indeed a subset of nums2. However, the LLM-generated precondition simply uses True, completely ignoring all stated constraints. This makes the precondition incomplete as it accepts invalid inputs that violate the problem’s fundamental assumptions, potentially leading to incorrect behavior in the implementation and proof generation phases. In our test suites, we have a negative test case with nums1 = [1, 1] and nums2 = [1, 2]. Since the LLM-generated pre-condition accepts this negative test case, our evaluation metric determines that the LLM-generated pre-condition is incomplete.